write equation of a perpendicular line in slope intercept form if y=2/5x-4 (-1,2)
Answer by nerdybill(7384)

write equation of a perpendicular line in slope intercept form if y=2/5x-4 (-1,2)
.
Slope of:
y=2/5x-4
is 2/5
.
A line perpendicular must have a slope that is the negative reciprocal:
(2/5)m = -1
m = -5/2 (this is the slope of the new line)
.
Plug slope (-5/2) and (1,2) into "point-slope" form:
y - y1 = m(x - x1)
y - 2 = (-5/2)(x - 1)
y - 2 = (-5/2)x + 5/2
y = (-5/2)x + 5/2 + 2
y = (-5/2)x + 5/2 + 4/2
y = (-5/2)x + 9/2 (this is what they're looking for)